# MoSE: Hierarchical Self-Distillation Enhances Early Layer Embeddings

### v2

@@ 5 SynthCoNL
We built SynthCoNL, a dataset that supports training and evaluation of encoders in text-to-code and code-to-code search, under multiple languages. Using the popular CodeSearchNet (Husain et al., 2019) as a seed dataset and selecting popular programming languages (Python, Java, Go, and PHP), we augmented it by transpiling available code snippets onto other languages.
To generate semantically similar code snippets for code-to-code search, we translated each snippet into a different language randomly sampled from Go, Ruby, Python, Java, C++, PHP, C, JavaScript. We prompted the Qwen2.5-Coder-7B-Instruct model with the source code, the name of the source language, and the name of the target language (see fig. 3). During code translation, we carried out a greedy sampling so as to prevent semantic discrepancies.
This process yielded pairs of code snippets in distinct languages tied to the same natural language description. As a result, every sample in the fine-tuning dataset includes a natural language description and two code snippets from distinct languages. SynthCoNL contains 1,071,367 triplets where, in the first code column, we directly sampled code snippets from CodeSearchNet, including Python, Java, PHP, and Go. The second code column, artificially synthesized via code translation, includes Go, Ruby, JavaScript, Python, C++, PHP, C, and Java code snippets. After a manual inspection, we discovered that both columns contained code snippets that differed only in identifiers or function arguments. Several tasks were semantically identical but paraphrased with different parameter requirements (e.g., two identical paraphrased tasks asked for opening a socket on a different port). During the preprocessing phase of SynthCoNL, motivated by the dataset’s redundancy and preliminary experiments that show its effectiveness on the model’s performance, we near-deduplicated the dataset using both the CodeSearchNet code column and the synthesized code column. During the data near deduplication phase, we relied on Locality Sensitive Hashing (LSH) with a Jaccard similarity threshold of 0.7 and 256 permutations, analyzing character-level 5-grams.
@@ 6.1 Evaluation
Figure 4: Performance of different models on CT and POJ104 for code-to-code retrieval with CodeXGLUE dataset. MoSE has state-of-the-art results with the CT benchmark and outperforms all the open-source models over the POJ104 benchmark. Compared to closed-source solutions (OpenAItext-embedding-3-large), MoSE performs on par in the code translation task and strongly reduces the gap between open- and closed-source models with the POJ104 benchmark.
We evaluated MoSE, after fine-tuned for retrieval, on diverse code understanding tasks, leveraging the CodeXGLUE (Lu et al., 2021) group of benchmarks.
@@
Additionally, MoSE was fine-tuned and evaluated on the separate BigCloneBench (Svajlenko & Roy, 2015) dataset for Java code clone detection (a binary classification task), assessing its classification performance.
@@ 6.2 Benchmarks
Table 1 presents the results for the CodeSearchNet (text-to-code) task in terms of Mean Reciprocal Rank (MRR, higher is better) for each single language, average Normalized Discounted Cumulative Gain (NDCG, higher is better), and average MRR. Results for UniXcoder, ModernBERT, and CodeT5+ are reported from the original papers (Guo et al., 2022; Warner et al., 2024; Wang et al., 2023). On CodeSearchNet, MoSE achieves an MRR of 81.0 and a NDCG of 84.2, outperforming CodeT5+ (770M), UniXcoder, and ModernBERT-large. The only encoder that surpasses MoSE is OpenAI’s text-embedding-3-large.
In Figure 4, we present results from both POJ104 and CT datasets reported respectively in terms of MRR for code translation (Java to C# retrieval) and mean average precision for POJ104 (C++ to C++ retrieval). MoSE reaches the best performance among the evaluated models. We further replicated the benchmarking for all models in a zero-shot setting for code-to-code tasks as our model does not integrate POJ104 and the code translation datasets in the training set. We present a comprehensive and detailed analysis of code retrieval errors in appendix C.
Figure 5: Similarity score Heatmap presenting results from permutation tests (10,000 permutations with $\alpha$ = 0.05) across different exit points for Python-to-Java retrieval. * indicates p < 0.05 and ** indicates p < 0.001. Different layers, despite a common training objective, yield different similarity scores.
Referring to Figure 4, on the POJ104 dataset in zero-shot, MoSE achieves a mean Average Precision (mAP, higher is better) of 75.9, which is state-of-the-art among open-sourced models. However, it is significantly behind OpenAI text-embedding-3-large. We underscore that a direct comparison with OpenAI text-embedding-3-large remains challenging because it is closed-source, and details such as model size, training methodology, or potential data contamination are undisclosed.
We observe a marked difference in performance depending on whether the task is framed as text-to-code or code-to-code retrieval. As shown in Figure 1(b), shallow layers yield better performance in the code-to-code setting, whereas for text-to-code retrieval, optimal performance emerges at layer 18.
@@
To further investigate this phenomenon, we compared similarity scores across layers to assess how each layer represents code snippets differently. Figure 5 shows permutation test results for a Python-to-Java retrieval task, highlighting significant variations in similarity scores that support our previous findings. A detailed analysis of these results is provided in appendix C.
@@
Our method remains competitive compared to state-of-the-art models. While CodeT5+ (770M) and UniXcoder achieve peaks F1 of 95.1 and 95.2, MoSE delivers on-par results (-0.9) by Layer 18, with the added benefit of a modular architecture.
@@ 6.3 Ablation Study
We conducted an ablation study for the Multi-Layer loss, by fine-tuning (starting from MoSE, pre-trained) on SynthCoNL, each exit point singularly (e.g., for the baseline on layer 18, we retain only the first 18 layers and fine-tune the model using just the projection head on that layer). Finally, we compared the baseline models with the corresponding results (Self-Distilled) of MoSE fine-tuned for retrieval with the multi-layer loss; all tests are conducted on the CodeSearchNet dataset. MoSE consistently outperforms the single-exit baselines (up to +4.36% averaged Recall@1 gain on layer 9), indicating that lower-level layers benefit from training signals propagated from deeper layers. This behavior is highlighted in Table 3, where MoSE, indicated as Self-Distilled, outperforms all the single exit baselines consistently. This finding underscores a promising new direction in Self-Distillation for large-scale code and text models, enabling high performance even in more compact configurations. Moreover, Figure 6 illustrates that MoSE maintains robust performance from layers 18 to 36, allowing users to scale down the network to match their memory, computational, or latency constraints while preserving strong retrieval performances.
Table 3: Performance comparison of MoSE layers and baseline fine-tuned models on the CodeSearchNet benchmark. The table displays the overall retrieval performance measured by the Mean Reciprocal Rank (MRR, higher is better). We refer to MoSE, fine-tuned with multiple exit points simultaneously, as Self-Distilled. The models not marked as self-distilled are the baselines, fine-tuned individually for each exit point. MoSE fine-tuned with multi-layer loss (Self-Distillation) consistently outperforms the baselines, showing a more significant difference in the initial layers.
CodeSearchNet
Model type Num. Layers Num. Params. Ruby Javascript Go Python Java PHP avg. MRR
Base 4 $160M$ 59.5 61.3 72.1 86.2 68.2 75.5 70.5
Self-Distilled 62.2 64.7 74.8 88.1 71.4 78.0 73.2
Base 9 $300M$ 64.9 65.7 74.3 87.3 72.0 78.8 73.8
Self-Distilled 67.6 69.4 78.9 90.2 75.5 82.3 77.3
Base 18 $550M$ 73.8 73.5 82.4 92.1 78.4 84.0 80.7
Self-Distilled 74.1 74.0 82.5 92.5 78.7 84.5 81.0
Base 27 $800M$ 72.3 71.8 80.8 90.8 76.9 82.3 79.1
Self-Distilled 73.2 73.3 81.7 92.1 77.8 83.8 80.3
Base 36 $1B$ 72.3 72.9 80.7 91.5 77.1 82.9 79.5
Self-Distilled 73.5 72.6 80.5 91.4 76.9 82.7 79.6
@@ 7 Related work
Figure 6: Performance Comparison between MoSE fine-tuned with Multi-layer loss (Self-Distilled model) and baselines fine-tuned just on single exit points: The graph illustrates averaged MRR and Recall@1 results for different layers over the CodeSearchNet test set. Layers fine-tuned with Multi-Layer loss outperforms the baselines (up to 3.47% MRR and +4.36% Recall@1).
@@
Since the introduction of ELMo (Peters et al., 2018), deep contextual information has enhanced generating embeddings for textual retrieval or classification, reaching state-of-the-art results in several tasks. BERT (Feng et al., 2020) followed those findings, adapting the Transformer architecture (Vaswani et al., 2017) to enable a bi-directional representation with two different training objecting, namely the masked language modeling and the next sentence prediction losses. Lan et al. (2019) and Liu et al. (2019) adapted the BERT architecture to obtain an enhanced pre-trained model by removing or modifying the NSP, focusing on pre-training data or hyperparameters optimization. More recently, modernBERT (Warner et al., 2024) tied the gap between modern decoders (Jiang et al., 2023; Hui et al., 2024; Dubey et al., 2024; Touvron et al., 2023; Lozhkov et al., 2024) advancements that rely upon models with an increased number of parameters, trained upon more tokens, and being capable of handling longer contextual information.
@@
In code representation, large language models must be adapted by training them on a curated corpus focused on software and by leveraging code’s syntactic and semantic structures, which differ significantly from natural language. Feng et al. (2020) adapted the BERT architecture to produce semantically meaningful embeddings for source code, resulting in codeBERT. This was accomplished by including more source code in the training set and focusing on a training loss that can leverage bimodal (natural language and code) contextual information (Clark et al., 2020). GraphCodeBERT enhanced codeBERT (Feng et al., 2020) representations by incorporating data flow graphs, capturing dependencies between variables and operations, and improving tasks like code summarization and clone detection. UniXcoder (Guo et al., 2022) extended this by introducing a unified encoder-decoder framework, integrating abstract syntax trees (ASTs) and data flow information. Wang et al. (2023) expanded these findings with CodeT5+, stressing how multiple losses that leverage code semantics impact the model pertaining. The work incorporated text-code contrastive learning, text-code Matching, and text-code causal LM for better code understanding and generation.
When trying to achieve better performance, research has shifted toward models with a high number of parameters. While this trend appears effective from a performance perspective, end users may face computational or memory limitations as LLMs vary from millions to billions of parameters. Sanh et al. (2019) pioneered the introduction of knowledge distillation, using a “teacher” model that guides a smaller model to emulate its behavior. This methodology has been widely adopted and improved upon recently (DeepSeek-AI et al., 2025; Hui et al., 2024), becoming a standard for obtaining high-performing smaller LLMs.
@@
Our work differs from previous work by adapting a modern architecture (StarCoder-2 (Lozhkov et al., 2024)) to a code encoder-only based model and introducing a novel ’Self-Distillation’ mechanism. We replace the next sentence prediction loss with an In-Context Classification focused on the repository level and expand the context to 2,048 tokens. Our novel Self-Distillation mechanism improves low-level layers, resulting in a modular transformer architecture without additional teacher models or further data for distillation.
@@ 8 Conclusion
We introduced MoSE, a modular, multi-exit encoder that uses Self-Distillation to enhance early-layer representations while achieving state-of-the-art results on code understanding benchmarks. By integrating multiple exit points, MoSE enables explicit efficiency-accuracy trade-offs tailored to deployment constraints. Our approach not only allows for modular use but also enhances the alignment between task requirements and representation depth. This improves task-specific layer specialization, providing the end-user with a more adaptable model.
Future work will explore combining exit-points and further understanding task type and depth interactions. While our results are encouraging, we acknowledge that computational constraints limited extensive hyperparameter tuning and pre-training ablations, and that fine-tuning with synthetic code generated from code translation has an unclear impact. We discuss these limitations in detail in appendix A.
@@ Acknowledgments
We acknowledge ISCRA for awarding this project access to the LEONARDO supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CINECA (Italy).
This work was supported by Future AI Research (FAIR) PE01, SPOKE 8 on PERVASIVE AI funded by the National Recovery and Resilience Plan (NRRP).
Appendix A Limitations
Due to our dependence on multiple GPUs, we encountered significant computational constraints. Parameter grid searches using smaller and earlier models were the only methods to extrapolate the best hyperparameter setup. The optimal hyperparameters for smaller models can differ from those for larger ones; consequently, we encountered a limitation in discovering an effective training setup. Conducting an ablation study for the in-context classification loss was not practicable due to the high costs of the pre-training stage. Therefore, computational resources pose a significant constraint in this work, and we want to emphasize how this factor can also undermine the possibility of replicating the experiments.
Here, we highlight potential threats to the validity of the research process, focusing on both external and internal factors.
External validity
When synthesizing the SynthCoNL code, we rely on code translation; we understand that synthesized data adheres to stylistic writing patterns distinct from those of humans. We tested the model’s performance on standard benchmarks. However, the impact of utilizing code snippets as synthetic data in training large language models for generalization over human text-to-code and code-to-code search is still not fully understood.
Internal validity
The ablation study focused on fine-tuning the model with and without multi-layer loss. However, this comparison does not account for how the model behaves when starting from a model not pre-trained on multi-layer loss. Although our experiments present promising results, further inspection is necessary to better understand this phenomenon.
Appendix B Synthetic dataset
Table 4: SynthCoNL composition: Average character count and sample size per language for the CodeSearchNet seed dataset and synthesized portion obtained through translation.
Language CSN samples CSN avg. char Synth. samples Synth. avg. char
English 1,071,367 180 - -
PHP 280,706 514 116,967 579
Python 274,454 474 117,374 518
Go 234,089 350 124,125 541
Java 282,118 505 116,098 707
C++ - - 141,956 938
Ruby - - 158,494 456
C - - 136,365 1,029
JavaScript - - 159,988 557
SynthCoNL is a fine-tuning dataset designed for text-to-code and code-to-code search, built by augmenting CodeSearchNet Husain et al. (2019) with transpiled code snippets across multiple languages (Python, Java, Go, PHP, Ruby, C++, C, JavaScript). The dataset underwent a preprocessing phase, including deduplication based on the original and synthesized code columns. Near-deduplication was performed using Locality Sensitive Hashing (LSH) with a Jaccard similarity threshold of 0.7 over character-level 5-grams to remove semantically identical snippets differing only in identifiers or function arguments.
For code-to-code search, we translated each snippet into a randomly sampled target language using the Qwen2.5-Coder-7B-Instruct model with greedy search to ensure consistency. Each dataset entry consists of a natural language description and two code snippets in different languages. SynthCoNL contains 1,071,367 samples, with original code from CodeSearchNet (Python, Java, PHP, Go) and translated code (Go, Ruby, JavaScript, Python, C++, PHP, C, Java). In Figure 11, In Figure 12 and Figure 10 some examples of code translation are shown.
Table 4 shows the average number of characters per language in SynthCoNL, we emphasize that synthesized data is significantly longer than human-written code and might have stylistic differences compared to human code, we further discussed this in the Limitations section.
Appendix C Failure Case Analysis in Code-to-Code Retrieval
Figure 7: Retrieved incorrect samples from CodeT5+, UniXcoder, and the 18th layer of MoSE. All the models appear to be misled by structural similarities, such as loop functions and return values, highlighting a common superficial retrieval phase.
Figure 8: Histograms of normalized similarity score between correct and incorrect retrieved code snippets. The first 5 distributions belong respectively to MoSE 4, 9, 18, 27, 36 layers’ exit points; CodeT5+ and UniXcoder are the last ones displayed. While our model tends to have different, distinct distributions regarding wrong and correct retrievals, both UniXcoder and CodeT5+ have more overlapping results.
Table 5: Permutation tests between correct and incorrect similarity scores ($\alpha=0.05$). The table shows a big discrepancy between MoSE t-statistics and both UniXcoder and CodeT5+ results.
Model t-statistic p-value
MoSE (L4) 8.0 <0.001
MoSE (L9) 7.3 <0.001
MoSE (L18) 8.0 <0.001
MoSE (L27) 7.8 <0.001
MoSE (L36) 6.3 <0.001
UniXcoder 2.2 0.04
CodeT5+ 110M-embedding 2.7 0.009
Figure 9: Heatmap of permutation tests (10,000 permutations) between MoSE similarity score results ($\alpha=0.05$). * with p<0.05 and ** with p<0.001. In the left heatmap, we present the differences in correctly retrieved snippets across various layers, while the right heatmap illustrates the incorrectly retrieved snippets. The heatmaps demonstrate that even layers trained to achieve similar representation objectives can yield different similarity scores.
In this section, we analyzed where our model fails in code-to-code retrieval, relying upon gfg, a Python-to-Java dataset. We compared how our model gets misled by different code snippets with CodeT5+ and UniXcoder, and the confidence level with which the model identifies the occurrences. In fig. 7 we highlight a common defect of the most used open code encoders. When retrieving code snippets focusing on semantic similarity, we emphasize how all the inspected models might get misled by structural factors such as loops, if statements, and return types. The retrieved results underline an overconsideration of structural features despite semantic differences.
We then analyzed whether the model’s confidence might suggest uncertainty in retrieving incorrect snippets of code, indicating a statistically significant level of "awareness" in overconsidering misleading features. Thus, we compared correct and incorrect similarity distribution scores for retrieved code snippets. Figure 8 shows how MoSE, CodeT5+, and UniXcoder results differ when comparing correct and incorrect retrievals. In table 5, MoSE demonstrated a more significant t-statistic than CodeT5+ and UniXcoder. In a retrieval use case scenario, when the confidence level is insufficient, we suggest that thresholding the similarity score could be an option when deciding whether to provide the end user with a single result or a herd of results.
In fig. 9, we present the differences in similarity scores among the five exit points using a permutation test. It is important to note that, although the model is trained to maintain similar retrieval capabilities across these five exit points (as discussed in section 2.1), and while the incorrectly retrieved snippets might overlap between the layers, each exit point encodes code snippets differently.
Based on these results, we strongly recommend conducting a grid search between the similarity score threshold and the exit point to utilize.
Figure 10: Example of a code snippet translation from Python to PHP.
Figure 11: Example of a code snippet translation from Go to Ruby.
Figure 12: Example of a code snippet translation from Java to C.
Appendix D Model card
Table 6 provides a comprehensive overview of the hyperparameters, architectural choices, and training configurations used in the development of MoSE. This includes detailed settings for both the pre-training and fine-tuning stages, as well as hardware specifications. The information is provided to enhance reproducibility and improve understanding of the model’s design.
Table 6: Details on hyperparameters, architecture, and training for MoSE are presented here, including insights from both pre-training and fine-tuning stages.
Architecture Pre-training
Hyperparameter Value Hyperparameter Value
Model size 1B parameters Batch size 3.99M tokens
Precision bfloat16 Pretraining steps 245,000
Hidden layers 36 Pretraining tokens 1T
Attention heads 16 Loss function MLM + In-Context loss
Hidden dimensionality 1,024 Multi-layer loss Yes
Positional encoding RoPE ($\theta=10^{-6}$) Optimizer AdamW
Context length 2,048 Weight decay $1\mathrm{e}{-1}$
Attention mechanism Grouped Query Attention Initial learning rate $6.24\mathrm{e}{-4}$
Attention pattern Bi-directional Learning rate schedule Multi-step
Warmup steps 4,000
Fine-tuning Hardware (Pre-training + Fine-tuning)
Hyperparameter Value Hyperparameter Value
Dataset size 635,404 samples Pre-training GPUs 512 NVIDIA Ampere A100 (64GB)
Fine-tuning steps 20,000 Fine-tuning GPUs 512 NVIDIA Ampere A100 (64GB)
Loss function CLIP loss Benchmarking GPUs 1 NVIDIA Ampere A100 (64GB)
Multi-layer loss Yes Overall training hours 450,000
Batch size 2,048
Learning rate $1.0\mathrm{e}{-5}$
Temperature parameter 10.0

### v3

@@ 5 SynthCoNL
We built SynthCoNL, a dataset that supports training and evaluation of encoders in text-to-code and code-to-code search, under multiple languages. Using the popular CodeSearchNet (Husain et al. 2019) as a seed dataset and selecting popular programming languages (Python, Java, Go, and PHP), we augmented it by transpiling available code snippets onto other languages.
To generate semantically similar code snippets for code-to-code search, we translated each snippet into a different language randomly sampled from Go, Ruby, Python, Java, C++, PHP, C, JavaScript. We prompted the Qwen2.5-Coder-7B-Instruct model with the source code, the name of the source language, and the name of the target language (see Figure 3). During code translation, we carried out a greedy sampling so as to prevent semantic discrepancies. This process yielded pairs of code snippets in distinct languages tied to the same natural language description. As a result, every sample in the fine-tuning dataset includes a natural language description and two code snippets from distinct languages. SynthCoNL contains 1,071,367 triplets where, in the first code column, we directly sampled code snippets from CodeSearchNet, including Python, Java, PHP, and Go. The second code column, artificially synthesized via code translation, includes Go, Ruby, JavaScript, Python, C++, PHP, C, and Java code snippets. After a manual inspection, we discovered that both columns contained code snippets that differed only in identifiers or function arguments. Several tasks were semantically identical but paraphrased with different parameter requirements (e.g., two identical paraphrased tasks asked for opening a socket on a different port). During the preprocessing phase of SynthCoNL, motivated by the dataset’s redundancy and preliminary experiments that show its effectiveness on the model’s performance, we near-deduplicated the dataset using both the CodeSearchNet code column and the synthesized code column. During the data near deduplication phase, we relied on Locality Sensitive Hashing (LSH) with a Jaccard similarity threshold of 0.7 and 256 permutations, analyzing character-level 5-grams.
@@ 6.1 Evaluation
We evaluated MoSE, after fine-tuned for retrieval, on diverse code understanding tasks, leveraging the CodeXGLUE (Lu et al. 2021) group of benchmarks.
@@
Additionally, MoSE was fine-tuned and evaluated on the separate BigCloneBench (Svajlenko and Roy 2015) dataset for Java code clone detection (a binary classification task), assessing its classification performance.
@@ 6.2 Benchmarks
Table 1 presents the results for the CodeSearchNet (text-to-code) task in terms of Mean Reciprocal Rank (MRR, higher is better) for each single language, average Normalized Discounted Cumulative Gain (NDCG, higher is better), and average MRR. Results for UniXcoder, ModernBERT, and CodeT5+ are reported from the original papers (Guo et al. 2022; Warner et al. 2024; Wang et al. 2023). On CodeSearchNet, MoSE achieves an MRR of 81.0 and a NDCG of 84.2, outperforming CodeT5+ (770M), UniXcoder, and ModernBERT‐large. The only encoder that surpasses MoSE is OpenAI’s text-embedding-3-large.
In Table 1, we also present results from both POJ104 and CT datasets (code-to-code) reported respectively in terms of MRR for code translation (Java to C# retrieval) and mean average precision for POJ104 (C++ to C++ retrieval). MoSE reaches the best performance among the evaluated models. We further replicated the benchmarking for all models in a zero-shot setting for code-to-code tasks, as our model does not integrate POJ104 and the code translation datasets in the training set. We present a comprehensive and detailed analysis of code retrieval errors in the Appendix.
Figure 4: Similarity score Heatmap presenting results from permutation tests (10,000 permutations with $\alpha$ = 0.05) across different exit points for Python-to-Java retrieval. * indicates p << 0.05 and ** indicates p << 0.001. Different layers, despite a common training objective, yield different similarity scores.
On the POJ104 dataset, in zero-shot, MoSE achieves a mean Average Precision (mAP, higher is better) of 75.9, which is the highest among open-source models to our knowledge. However, it is significantly behind OpenAI text-embedding-3-large. We underscore that a direct comparison with OpenAI text‐embedding‐3‐large remains challenging because it is closed‐source, and details such as model size, training methodology, or potential test data leakage are undisclosed. We observe a difference in performance profiles depending on whether the task is framed as text-to-code or code-to-code retrieval. As shown in Figure 1(b), shallow layers yield better performance in the code-to-code setting, whereas for text-to-code retrieval, optimal performance emerges at layer 18.
@@
To further investigate this phenomenon, we compared similarity scores across layers to assess how each layer represents code snippets differently. Figure 4 shows permutation test results for a Python-to-Java retrieval task, highlighting significant variations in similarity scores that support our previous findings. A detailed analysis of these results is provided in the Appendix.
@@
Our method remains competitive compared to open models. While CodeT5+ (770M) and UniXcoder achieve peaks F1 of 95.1 and 95.2, MoSE delivers on-par results (-0.9) by Layer 18, with the added benefit of a modular architecture.
@@ 6.3 Ablation Study
Effectiveness of the Multi Layer Loss
We conducted an ablation study for the multi-layer loss, by fine-tuning each exit point independently on SynthCoNL, starting from the pre-trained MoSE. For example, for the baseline on layer 18, we retain the first 18 layers and fine-tune the model using just the projection head on that layer. Finally, we compared the baseline models with the corresponding results (Self-Distilled) of MoSE fine-tuned for retrieval with the multi-layer loss. All tests are conducted on the CodeSearchNet dataset. MoSE consistently outperforms the single‐exit baselines (up to +4.36% averaged Recall@1 gain on layer 9), indicating that lower‐level layers benefit from training signals propagated from deeper layers. This behavior is highlighted in Figure 5 where MoSE, indicated as Self-Distilled, outperforms all the single exit baselines consistently. This finding underscores a promising new direction in Self‐Distillation for large‐scale code and text models, enabling high performance even in more compact configurations. Moreover, Figure 5 also illustrates that MoSE maintains robust performance from layers 18 to 36, allowing users to scale down the network to match their memory, computational, or latency constraints while preserving strong retrieval performances.
Effectiveness of the In-Context Classification Objective
CoIR (NDCG@10)
Model CSN‑CCR CoSQA CT‑DL CT‑Contest
MoSE (ICC) 10.1 0.4 32.0 12.1
MoSE (NSP) 6.2 0.4 31.0 12.0
Table 3: Ablation study comparing pre-training objectives (In-Context Classification vs. Next Sentence Prediction), evaluated with NDCG@10 (%) on the CoIR suite. ICC demonstrates gains on the cross-context CSN‑CCR task.
We conducted an ablation study, detailed in Table 3, to assess whether our In-Context Classification (ICC) loss promotes inter-language understanding. For this study, we pre-trained two small versions of the model (each with 30M parameters) on 3B tokens. We compared the ICC loss with the Next Sentence Prediction (NSP) loss, commonly used in many state-of-the-art encoders (Li et al. 2023; Feng et al. 2020a).
To assess the performance of our models, we utilized the inter-domain section of the CoIR evaluation suite (Li et al. 2025), which focuses on tasks across various languages and domains. Our results indicated that the In-Context Classification loss either outperformed or matched the effectiveness of the Next Sentence Prediction loss, while providing a more efficient and dense representation. These results suggest that ICC may serve as a stronger default pretraining signal than NSP for multilingual and cross-context information retrieval tasks.
@@ 7 Related Work
Figure 5: Performance Comparison between MoSE fine-tuned with multi-layer loss (self-distilled model) and baselines fine-tuned just on single exit points: The graph illustrates averaged MRR and Recall@1 results for different layers over the CodeSearchNet test set. Layers fine-tuned with Multi-Layer loss outperforms the baselines (up to +4.36% Recall@1).
@@
Since the introduction of ELMo (Peters et al. 2018), deep contextual information has enhanced generating embeddings for textual retrieval or classification, reaching state-of-the-art results in several tasks. BERT (Feng et al. 2020a) followed those findings, adapting the Transformer architecture (Vaswani et al. 2017) to enable a bi-directional representation with two different training objectives, namely the masked language modeling and the next sentence prediction losses. (Lan et al. 2019) and (Liu et al. 2019) adapted the BERT architecture to obtain an enhanced pre-trained model by removing or modifying the NSP, focusing on pre-training data or hyperparameters optimization. More recently, modernBERT (Warner et al. 2024) tied the gap between modern decoders (Jiang et al. 2023; Hui et al. 2024; Dubey et al. 2024; Touvron et al. 2023; Lozhkov et al. 2024) advancements that rely upon models with an increased number of parameters, trained upon more tokens, and being capable of handling longer contextual information.
@@
In code representation, large language models must be adapted by training them on a curated corpus focused on software and by leveraging code’s syntactic and semantic structures, which differ significantly from natural language. Feng et al. (2020b) adapted the BERT architecture to produce semantically meaningful embeddings for source code, resulting in codeBERT. This was accomplished by including more source code in the training set and focusing on a training loss that can leverage bimodal (natural language and code) contextual information (Clark et al. 2020). GraphCodeBERT enhanced codeBERT (Feng et al. 2020b) representations by incorporating data flow graphs, capturing dependencies between variables and operations, and improving tasks like code summarization and clone detection. UniXcoder (Guo et al. 2022) extended this by introducing a unified encoder-decoder framework, integrating abstract syntax trees (ASTs) and data flow information. Wang et al. (2023) expanded these findings with CodeT5+, stressing how multiple losses that leverage code semantics impact the model pertaining. The work incorporated text-code contrastive learning, text-code Matching, and text-code causal LM for better code understanding and generation. When trying to achieve better performance, research has shifted toward models with a high number of parameters. While this trend appears effective from a performance perspective, end users may face computational or memory limitations as LLMs vary from millions to billions of parameters. Sanh et al. (2019) pioneered the introduction of knowledge distillation, using a “teacher” model that guides a smaller model to emulate its behavior. This methodology has been widely adopted and improved upon recently (DeepSeek-AI et al. 2025; Hui et al. 2024), becoming a standard for obtaining high-performing smaller LLMs.
@@
Our work differs from previous work by adapting a modern architecture (StarCoder-2 (Lozhkov et al. 2024)) to a code encoder-only based model and introducing a novel ’Self-Distillation’ mechanism. We replace the next sentence prediction loss with an In-Context Classification focused on the repository level and expand the context to 2,048 tokens. Our novel Self-Distillation mechanism improves low-level layers, resulting in a modular transformer architecture without additional teacher models or further data for distillation.
@@ 8 Conclusion
We introduced MoSE, a modular, multi-exit encoder that uses Self-Distillation to enhance early-layer representations while achieving competitive results on code understanding benchmarks. By integrating multiple exit points, MoSE enables explicit efficiency-accuracy trade-offs tailored to deployment constraints. Our approach not only allows for modular use but also enhances the alignment between task requirements and representation depth. This improves task-specific layer specialization, providing the end-user with a more adaptable model. Future work will explore combining exit-points and further understanding task type and depth interactions. While our results are encouraging, we acknowledge that computational constraints limited extensive hyperparameter tuning and pre-training ablations, and that fine-tuning with synthetic code generated from code translation has an unclear impact.
@@ Acknowledgments
We acknowledge ISCRA for awarding this project access to the LEONARDO supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CINECA (Italy). This work was supported by Future AI Research (FAIR) PE01, SPOKE 8 on PERVASIVE AI funded by the National Recovery and Resilience Plan (NRRP).